A flywheel with a radius of 0.300 m starts from rest and accelerates with a constant angular acceleration of 0.200 rad/s2 .

(a) Compute the magnitude of the resultant acceleration of a point on its rim after it has turned through 60.0 degrees.

(b) Compute the magnitude of the radial acceleration of a point on its rim after it has turned through 120.0 degrees.
Answer #1

Radius of the flywheel = R = 0.3 m

Initial angular speed of the flywheel = 1 = 0 rad/s

Angular acceleration of the flywheel = = 0.2 rad/s2

Angle through which the flywheel has turned = 1 = 60 deg = 60 x (2/360) rad = 1.0472 rad

Angular speed of the flywheel after it has turned through 60 degrees = 2

22 = 12 + 21

22 = (0)2 + (2)(0.2)(1.0472)

2 = 0.6472 rad/s

Radial acceleration of a point at the rim after the flywheel has turned through 60 degrees = a1r

a1r = 22R

a1r = (0.6472)2(0.3)

a1r = 0.12566 m/s2

Tangential acceleration of a point at the rim after the flywheel has turned through 60 degrees = a1t

a1t = R

a1t = (0.2)(0.3)

a1t = 0.06 m/s2

Resultant acceleration of a point at the rim after the flywheel has turned through 60 degrees = a1

a1 = 0.139 m/s2

Angle through which the flywheel has turned = 2 = 120 deg = 120 x (2/360) rad = 2.0944 rad

Angular speed of the flywheel after it has turned through 120 degrees = 3

32 = 12 + 22

32 = (0)2 + (2)(0.2)(2.0944)

3 = 0.9153 rad/s

Radial acceleration of a point at the rim after the flywheel has turned through 120 degrees = a2r

a2r = 32R

a2r = (0.9153)2(0.3)

a2r = 0.251 m/s2

a) Magnitude of the resultant acceleration of a point on the rim after it has turned through 60 degrees = 0.139 m/s2

b) Magnitude of the radial acceleration of a point on the rim after it has turned through 120 degrees = 0.251 m/s2